OpenStudy (anonymous):

Use square root property to solve this equation, explanations are loved :) its: (z+1)^2=17

OpenStudy (slaaibak):

z = sqrt(17) - 1 z= -sqrt(17) - 1

OpenStudy (anonymous):

\[1-\sqrt{17}, 1+\sqrt{17}\] Is this the correct way to type it as an answer?

OpenStudy (anonymous):

\[(z+1)^{2} = 17\] You need to get rid of the squared so to do that you take square root on both sides..\[\sqrt{(z+1)^{2}} = \sqrt{17}\] The square root cancels out the squared leaving you with \[z + 1 = \sqrt{17}\] Subtract 1 from both sides leaving you with the final answer as \[z = \sqrt{17} - 1\]

OpenStudy (anonymous):

I forgot the plus minus sign

OpenStudy (anonymous):

do you write it as a negative then you put a comma then as positive? two answers?

OpenStudy (anonymous):

\[z = \pm \sqrt{17} - 1\]

OpenStudy (anonymous):

awesome! Thanks everyone :3